Solutions of iron (III) chloride, FeCl3, are used in photoengraving and to make ink. This compound can be made by the following reaction:



a.) How many grams of Fe are needed to combine with 4.5 moles of Cl2?

b.) If 240 g of Fe is to be used in this reaction, with adequate Cl2, how many moles of FeCl3 will be produced
1
Expert's answer
2013-01-15T08:24:05-0500

2Fe + 3Cl2 = 2FeCl3

n(Fe) = (4.5mol*3)/2 = 6.75 mol

n(Fe) = m/M

n(Fe) = 240g/56g/mol = 4.285 mol

n(FeCl3) = 4.285 mol

M(FeCl3) = 56 + 35.5*3 = 162.5 g/mol

m(FeCl3) = 162.5g/mol * 4.285 mol = 696 g